. In a manufacturing organisation, the distribution of wages was perfectly



normal and the number of workers employed in the organisation was 5000.



The mean wages of the workers were calculated at Rs. 800 and standard



deviation was worked out to be Rs. 200. On the basis of the information



estimate:



(i) The number of workers getting salary between Rs.700 and Rs.900.



(ii) Percentage of workers getting salary above Rs.1000



(iii) Percentage of workers getting salary below Rs.600

1
Expert's answer
2022-06-16T10:33:01-0400

(i)


"P(700<X<900)=P(Z<\dfrac{900-800}{200})"

"-P(Z\le\dfrac{700-800}{200})"

"=P(Z<0.5)-P(Z\le-0.5)"

"\approx0.69146-0.30854\approx0.3829"

"0.3829(500)=1915" workers


(ii)


"P(X>1000)=1-P(Z\le\dfrac{1000-800}{200})"

"=1-P(Z\le1)\approx0.1587"

"15.87" %


(iii)


"P(X<600)=P(Z<\dfrac{600-800}{200})"

"=P(Z<-1)\approx0.1587"

"15.87" %